40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 10 02:28:04 2013, Annalisa, Update, Green Locking, Beat notes lost! 
    Reply  Wed Jul 10 11:27:02 2013, manasa, Update, Green Locking, X arm beatnote found 
Message ID: 8819     Entry time: Wed Jul 10 02:28:04 2013     Reply to this: 8820
Author: Annalisa 
Type: Update 
Category: Green Locking 
Subject: Beat notes lost! 

[Manasa, Jenne, Annalisa]

I was going to find the beat note to start the cavity scan, but I couldn't.

These are the steps I followed:

  • locked the arm with IR to reduce the arm swinging
  • locked the green on the arm
  • started changing the green temperature setting the offset from the slow servo2 in the ALS. The PSL slow actuator ADJ was always set approximately to zero, and the PSL temperature was checked in order to set the auxiliary laser temperature where the beat was expected (as in the plot)

After spanning the temperature by approximately 4degC, we started be suspicious that I couldn't find the beat in the range of temperature where it was supposed to be found, and we started making several trials:

  • PD output disconnected from the beatbox and connected to the cable running to the Control Room
  • Checked that the cable going to the Control Room was working by sending a signal with the Marconi (the cable was working)
  • Put back the amplifier that had been previously removed
  • PD DC output checked with the oscilloscope
  • Spectrum analyzer connected to the PD output without passing trough the cable

The same trials were done also for the X arm, but we didn't succeed in finding the beat for the X neither.

 

ELOG V3.1.3-